OpenStudy (anonymous):

need help with a integral (sin^2(t)cos^4(t)), I can split it up into: ((1-cos(2t))/2)((1+cos(2t))/2)^2, but the solution then takes out the constants and make it so its 1/8 infront of the integral, how do you get 1/8? I can get 1/4 but not 1/8

OpenStudy (turingtest):

The way I would do it is to apply\[\sin^2t=1-\cos^2t\]them use the double angle formula\[\cos^2t=\frac12(1+\cos(2t))\]

hartnn (hartnn):

you wanted to know how we get 1/8 ? [(1-...)/2 ] [(1+...)/2]^2 = [1/2-.....] [1/4+ ...] = [1/8 +....]

OpenStudy (anonymous):